REVISION DATE: 03-Apr-2018 11:51:11
CONFIG 13: 02-Apr-2018 to present
In Line PSBS custom made by Mark, aligned on laser bench Pier 35, S/N 2. Test of temperature/humidity stability of Collet Fiber Connector. Output Fiber: Collet Fiber on one end, FC on the other; used to replace the output fiber on Config11. Collet marked "DO NOT TWIST"; fiber labeled with orange electrical tape. This is the same fiber as Config12. Input Fiber: Collet Fiber on one end, FC on the other; used to replace the input fiber on Config12. Collet marked "DO NOT TWIST"; fiber labeled with green electrical tape. We took out the PDBS and the fold mirror. PD2 is connected to the through channel and PD1 has a 9V battery. SIRCUS Sphere, Yuqin Zong's white LED source, monitor with Si PD and SEI. PD2 now on Gain 6 * E1 = Day 1: 02-Apr-2018 22degC, 22/30/22 at 30%RH with darks afterFigure 1: All data in this Configuration to date. PD signals are normalized to the monitor. Ratio PD1/Mon, Monitor, and normalized Signals all normalized to the initial value.
Figure 2: Zoomed in look at Day 1